In five years Dad will be three times as old as his daughter Jill will be then. If the sum of their present ages is 50, how old are they now?
Answer by Maths68(1474)

Let
Dad's current age = d
Jill's current age = j
Sum of their ages = 50
d+j=50
j=50-d......(1)
In five years means we have to add 5 in both ages
d+5=j+5
Now it says that then Dad's age will be 3 time of Jill's age
d+5=3(j+5)
d+5=3j+15
d=3j+15-5
d=3j+10..........(2)
Put the value of j in (2)
d=3(50-d)+10
d=150-3d+10
d=160-3d
d+3d=160
4d=160
4d/4=160/4
d=40
Put the value of d in (1)
j=50-d......(1)
j=50-40
j=10
So Jill is 10 and Dad is 40 years old